Find the Unit Normal Vector and Unit Binormal Vector:

( 1 point) Consider the helix r(t) (cos(8t), sin(8t),-3t). Compute, at- A, The unit tangent vector T-〈10.8 10884854070| , -0.46816458878| B. The unit normal vector N 〈 C. The unit binormal vector B-〈 1 ǐ ,1-0.35 11 23441 58 0
Answer #1

T(t) N(t) = F (t) Pr에

B(t) = T(t) x N(t)

N=(0.5,0.866,0)

B=(0.3,-0.175,0.936)